# High-Throughput and Scalable Secure Inference Protocols for Deep Learning with Packed Secret Sharing

### v1

@@ V-A Parallel Prefix-ORs and Bitwise Less Than
    We first describe our method to compute the prefix-ORs, which is the core and bottleneck to realizing the bitwise less-than functionality. The prefix-ORs operation ($b_{j}=\lor_{i=0}^{j}a_{i},a_{i}\in\{0,1\},j\in[\ell]$) can be implemented via its dual problem, that is, $\bar{b}_{j}=\land_{i=0}^{j}\bar{a}_{i}.$ It is worth noting that [5] designs a protocol for prefix-ORs with only one round of communication; however, [11] identified a privacy leakage issue in that protocol. In this work, we avoid this privacy leakage by utilizing functionality $\mathcal{F}_{Pmult-DN}$ and provide a secure protocol implementation for parallel prefix-ORs in Algorithm 8. $\mathcal{F}_{PreMult}$ denotes parallel prefix multiplication functionality and we formally define the functionality in Appendix -A. Further, $\mathcal{F}_{PreMult}$ can be securely implemented by directly employing $\mathcal{F}_{Pmult-DN}$ as a subroutine. As illustrated in Fig. 6, an 8-input prefix multiplication is depicted, where denotes a multiplication gate.
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil$ rounds with $\lceil\ell\text{log}_{2}(\ell)\rceil\frac{n}{k}$ field elements per party. The offline communication complexity is 1 rounds with $\frac{n\ell}{k(n-t)}\lceil\text{log}_{2}(\ell)\rceil$ field elements per party.
@@
    Theorem 8.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    Based on PSS, multiple bitwise less-than circuits can be securely computed in parallel. Specifically, given $k$ public values $a_{0},\cdots,a_{k-1}$ and $\ell$ shared bits vectors $\llbracket\boldsymbol{b}^{i}\rrbracket_{d},i\in[\ell],\boldsymbol{b}^{i}\in\{0,\cdots,1\}^{k}$, it outputs the share of $(a_{j}<b_{j}),b_{j}=\sum_{i=0}^{\ell-1}2^{i}\boldsymbol{b}^{i}_{j},j\in[k]$. Our protocol $\Pi_{Bitwise-LT}$ follows the method in [5], with the substitution of our secure prefix-ORs subprotocol $\Pi_{PreOR}$. Algorithm 9 depicts the protocol $\Pi_{Bitwise-LT}$. In Step 4, the XOR operation can be computed by $\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}\oplus\llbracket\boldsymbol{b}^{i}\rrbracket_{d}=\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}+\llbracket\boldsymbol{b}^{i}\rrbracket_{d}-2\cdot\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}\cdot\llbracket\boldsymbol{b}^{i}\rrbracket_{d}.$ We formally define the functionality $\mathcal{F}_{Xor}$ in Appendix -A and provide a secure protocol implementation in Appendix -B.
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil+2$ rounds with $\frac{n}{k}(\frac{4}{n-t}+\frac{1}{n}+\lceil\ell\text{log}_{2}(\ell)\rceil)$ field elements per party. The offline communication complexity is 1 rounds with $\frac{n}{k(n-t)}(4+\lceil\ell\text{log}_{2}(\ell)\rceil)$ field elements per party.
    Theorem 9.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    Figure 6: An 8-input prefix multiplication. denotes a multiplication gate that can be securely implemented by $\mathcal{F}_{Pmult-DN}$.
@@ V-B Parallel DReLU and ReLU
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil+5$ rounds with $\frac{n}{k}(\frac{4}{n-t}+\frac{1}{n}+\lceil\ell\text{log}_{2}(\ell)\rceil+4)+n$ field elements per party. The offline communication complexity is 4 rounds with $\frac{n}{k}(\frac{1}{n-t}(13+\lceil\ell\text{log}_{2}(\ell)\rceil)+\frac{4}{n}+k+1)$ field elements per party.
@@
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil+6$ rounds with $\frac{n}{k}(\frac{4}{n-t}+\frac{1}{n}+\lceil\ell\text{log}_{2}(\ell)\rceil+6)+n$ field elements per party. The offline communication complexity is 4 rounds with $\frac{n}{k}(\frac{1}{n-t}(15+\lceil\ell\text{log}_{2}(\ell)\rceil)+\frac{4}{n}+k+1)$ field elements per party.
@@
    Theorem 10.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
@@
      Output $\llbracket\text{ReLU}(\boldsymbol{a})\rrbracket_{d}\leftarrow\mathcal{F}_{Pmult-DN}(\llbracket\boldsymbol{a}\rrbracket_{d},\llbracket\boldsymbol{b}\rrbracket_{d}).$
@@
    Theorem 11.
    $\Pi_{ReLU}$ securely realizes $\mathcal{F}_{ReLU}$ in the $(\mathcal{F}_{DReLU},\mathcal{F}_{PMult-DN})$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
@@ V-C Parallel Maxpool
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(m)\rceil(\lceil\text{log}_{2}(\ell)\rceil+6)$ rounds with $(\frac{n}{k}(\frac{4}{n-t}+\frac{1}{n}+\lceil\ell\text{log}_{2}(\ell)\rceil+6)+n)(m-1)$ field elements per party. The offline communication complexity is 4 rounds with $\frac{n}{k}(\frac{1}{n-t}(15+\lceil\ell\text{log}_{2}(\ell)\rceil)+\frac{4}{n}+k+1)(m-1)$ field elements per party.
    TABLE I: Communication costs of per party of Liu et al. [5] and our protocol over different datasets and networks for secure inference. Communication is in MB. $n$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS, and OOM indicates memory overflow.
    11 3  [5] - 71.77 40.45 112.22 105.63 59.53 165.16 165.04 99.40 264.44 1975.53 1113.65 3089.18
    4  [5] - 67.81 43.17 110.98 99.81 63.54 163.35 154.42 106.07 260.49 1866.63 1188.51 3055.14
    21 3  [5] - 78.73 36.97 115.70 115.89 54.42 170.31 183.03 90.86 273.89 2167.38 1017.94 3185.32
    7  [5] - 72.29 42.93 115.22 106.40 63.20 169.60 169.30 105.50 270.80 1989.81 1182.10 3171.91
    31 3  [5] - 80.89 35.62 116.51 119.07 52.42 171.49 188.64 87.54 276.18 2226.84 980.78 3207.62
    12  [5] - 70.30 44.85 115.15 103.47 66.02 169.49 159.74 110.21 269.95 OOM OOM OOM
    63 3  [5] - 83.06 34.11 117.17 122.26 50.21 172.47 194.29 83.85 278.14 OOM OOM OOM
    20  [5] - 75.26 42.84 118.10 110.78 63.05 173.83 172.48 105.27 277.75 OOM OOM OOM
@@
    Theorem 12.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
@@ VI Evaluation
    TABLE II: Running time of Liu et al. [5] and our protocol over different datasets and networks for secure inference in WAN. Running time is in seconds. $n$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS, and OOM indicates memory overflow.
@@
    11 3  [5] - 17.30 9.64 26.94 23.80 13.92 37.72 35.35 22.15 57.50 426.93 182.71 609.64
    4  [5] - 16.82 9.93 26.75 23.51 13.21 36.72 34.34 22.43 56.77 420.59 183.26 603.85
    21 3  [5] - 32.82 15.68 48.50 48.14 22.02 70.16 64.87 35.80 100.67 914.20 307.68 1221.88
    7  [5] - 30.16 15.72 45.88 44.11 22.05 66.16 62.81 36.66 99.47 910.92 357.59 1268.51
    31 3  [5] - 48.14 20.09 68.23 67.59 28.47 96.06 95.06 46.27 141.33 1203.35 461.91 1665.26
    12  [5] - 44.12 22.01 66.13 63.73 30.32 94.05 90.15 49.95 140.10 OOM OOM OOM
    63 3  [5] - 89.99 36.69 126.68 131.76 55.08 186.84 188.09 86.27 274.36 OOM OOM OOM
    20  [5] - 87.73 45.19 132.92 124.20 64.03 188.23 180.91 106.59 287.50 OOM OOM OOM
    TABLE III: Running time of Liu et al. [5] and our protocol over different datasets and networks for secure inference in LAN. Running time is in seconds. $n$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS, and OOM indicates memory overflow.
@@
    In this section, we evaluate the performance of secure inference on 4 networks using MNIST and CIFAR10, and analyse the experimental results.
    VI-A Experiments Environment
    We implement this framework on top of Liu et al. [5] in C++ [^1] The source code will be open-sourced soon in github. and run our experiments on 11 Alibaba Cloud servers, each of which was equipped with an AMD EPYC 32-core processor with 64GB of RAM. To simulate different network conditions, we use Linux tc tool to simulate a wide-area network (WAN, RTT: 40 ms, 100 Mbps), and a local-area network (LAN, RTT: 0.2 ms, 10 Gbps). RTT denotes Round-Trip Time. We completed performance evaluations for inference in various settings involving different numbers of parties, including 11-party computation (11PC), 21PC, 31PC, and 63PC. In addition, we rerun the protocol of Liu et al. [5] under the same adversarial model. Since the single-round communication prefix multiplication protocol in  [5] suffers from privacy leakage issues, we replace it with a logarithmic-round prefix multiplication protocol based on the DN multiplication protocol. In the experiment, all arithmetic operations on the shares are performed modulo $p=2^{61}-1$, with a fixed-point precision of 13 bits. All evaluations are executed 10 times, and record the average.
    VI-B Neural Networks & Datasets

### v2

@@ V-A Parallel Prefix-ORs and Bitwise Less Than
    We first describe our method to compute the prefix-ORs, which is the core and bottleneck to realizing the bitwise less-than functionality. The prefix-ORs operation ($b_{j}=\lor_{i=0}^{j}a_{i},a_{i}\in\{0,1\},j\in[\ell]$) can be implemented via its dual problem, that is, $\bar{b}_{j}=\land_{i=0}^{j}\bar{a}_{i}.$ It is worth noting that LXY24 designs a protocol for prefix-ORs with only one round of communication; however, [31] identified a privacy leakage issue in that protocol. In this work, we avoid this privacy leakage by utilizing functionality $\mathcal{F}_{PDN}$ and provide a secure protocol implementation for parallel prefix-ORs in Algorithm 8. $\mathcal{F}_{PreMult}$ denotes parallel prefix multiplication functionality and we formally define the functionality in Appendix -C. Further, $\mathcal{F}_{PreMult}$ can be securely implemented by directly employing $\mathcal{F}_{PDN}$ as a subroutine. As illustrated in Fig. 6, an 8-input prefix multiplication is depicted, where denotes a multiplication gate.
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil$ rounds with $\frac{\lceil\ell\text{log}_{2}(\ell)\rceil}{k}$ field elements amortized per party. The offline communication complexity is 1 rounds with $\frac{n\ell}{k(n-t)}\lceil\text{log}_{2}(\ell)\rceil$ field elements amortized per party.
@@
    Based on PSS, multiple bitwise less-than circuits can be securely computed in parallel. Specifically, given $k$ public values $a_{0},\cdots,a_{k-1}$ and $\ell$ shared bits vectors $\llbracket\boldsymbol{b}^{i}\rrbracket_{d},i\in[\ell],\boldsymbol{b}^{i}\in\{0,\cdots,1\}^{k}$, it outputs the share of $(a_{j}<b_{j}),b_{j}=\sum_{i=0}^{\ell-1}2^{i}\boldsymbol{b}^{i}_{j},j\in[k]$. Our protocol $\Pi_{Bitwise-LT}$ follows the method in LXY24, with the substitution of our secure prefix-ORs subprotocol $\Pi_{PreOR}$. Algorithm 9 depicts the protocol $\Pi_{Bitwise-LT}$. In Step 4, the XOR operation can be computed by $\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}\oplus\llbracket\boldsymbol{b}^{i}\rrbracket_{d}=\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}+\llbracket\boldsymbol{b}^{i}\rrbracket_{d}-2\cdot\llbracket\boldsymbol{a}^{i}\rrbracket_{k-1}\cdot\llbracket\boldsymbol{b}^{i}\rrbracket_{d}.$ We formally define the functionality $\mathcal{F}_{Xor}$ in Appendix -C and provide a secure protocol implementation in Appendix -B.
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil+2$ rounds with $\frac{2\ell+2+\lceil\ell\text{log}_{2}(\ell)\rceil}{k}$ field elements amortized per party. The offline communication complexity is 1 rounds with $\frac{n}{k(n-t)}(4+\lceil\ell\text{log}_{2}(\ell)\rceil)$ field elements amortized per party.
    Figure 6: An 8-input prefix multiplication. denotes a multiplication gate that can be securely implemented by $\mathcal{F}_{PDN}$.
@@ V-B Parallel DReLU and ReLU
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil+5$ rounds with $\frac{\ell\mathrm{log}(\ell)+2\ell+7+k}{k}$ field elements amortized per party. The offline communication complexity is 4 rounds with $\frac{n}{k}(\frac{1}{n-t}(13+\lceil\ell\text{log}_{2}(\ell)\rceil)+\frac{4}{n}+k+1)$ field elements amortized per party.
@@
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(\ell)\rceil+6$ rounds with $\frac{\ell\mathrm{log}(\ell)+2\ell+9+k}{k}$ field elements amortized per party. The offline communication complexity is 4 rounds with $\frac{n}{k}(\frac{1}{n-t}(15+\lceil\ell\text{log}_{2}(\ell)\rceil)+\frac{4}{n}+k+1)$ field elements amortized per party.
@@
      Output $\llbracket\text{ReLU}(\boldsymbol{a})\rrbracket_{d}\leftarrow\mathcal{F}_{PDN}(\llbracket\boldsymbol{a}\rrbracket_{d},\llbracket\boldsymbol{b}\rrbracket_{d}).$
@@ V-C Parallel Maxpool
    To sum up, the online communication complexity is $\lceil\text{log}_{2}(m)\rceil(\lceil\text{log}_{2}(\ell)\rceil+6)$ rounds with $\frac{(m-1)(\ell\mathrm{log}(\ell)+2\ell+9+k)}{k}$ field elements amortized per party. The offline communication complexity is 4 rounds with $\frac{n}{k}(\frac{1}{n-t}(15+\lceil\ell\text{log}_{2}(\ell)\rceil)+\frac{4}{n}+k+1)(m-1)$ field elements amortized per party.
@@ VI Evaluation
    VI-A Experiments Environment
    We implement this framework on top of LXY24 in C++ and run our experiments on 11 Alibaba Cloud servers, each of which was equipped with an AMD EPYC 32-core processor with 64GB of RAM. We use the Eigen library to accelerate matrix multiplication and parallelize the computations using multithreading. To simulate different network conditions, we use Linux Traffic Control (tc) to simulate a wide-area network (WAN, RTT: 40 ms, 100 Mbps), and a local-area network (LAN, RTT: 0.2 ms, 10 Gbps). RTT denotes Round-Trip Time. We completed performance evaluations for inference in various settings involving different numbers of parties, including 11-party computation (11PC), 21PC, 31PC, and 63PC. In addition, we rerun the protocol of LXY24 under the same adversarial model. Since the single-round communication prefix multiplication protocol in LXY24 suffers from privacy leakage issues, we replace it with a logarithmic-round prefix multiplication protocol based on the DN multiplication protocol. In the experiment, all arithmetic operations on the shares are performed modulo $p=2^{61}-1$, with a fixed-point precision of 13 bits. All evaluations are executed 10 times, and the average is recorded.
    TABLE I: Online round and communication complexity of the main protocols. Numbers of the amortized communication are reported in field elements per party. The shapes of the fixed matrix multiplication with truncation are $u\times v$ and $v\times o$. 2D convolution $(\mathrm{Conv}_{m,c,o,f,s})$ with the inputs of size $m\times m$, $c$ input channels, $o$ kernels of size $f\times f$ , and the stride $s$. InLXY24, the input to the prefix-OR $(\mathrm{PreOR}_{\ell})$ is a single vector of length $\ell$. In ours, the input consists of $k$ vectors, each of length $\ell$. The input to several building blocks, including bitwise less than $(\mathrm{BitwiseLT}_{k})$, $\mathrm{DReLU}_{k}$, $\mathrm{ReLU}_{k}$ and $\mathrm{Maxpool}_{k}$, is a vector of length $k$.
    Protocol Rounds Communication
    LXY24 ours LXY24 ours
    $\mathrm{MatMult}_{u,v,o}(\Pi_{VecMatMult}^{Fixed})$ 1 1 $2uv$ $(k+1)uv/k$
    $\mathrm{Conv}_{m,c,o,f,s}(\Pi_{PMatMult}^{Fixed})$ 1 1 $2((m-f)/s+1)^{2}o$ $2((m-f)/s+1)^{2}o/k$
    $\mathrm{PreOR}_{\ell}(\Pi_{PreOR})$ $\mathrm{log}(\ell)$ $\mathrm{log}(\ell)$ $\ell\mathrm{log}(\ell)$ $\ell\mathrm{log}(\ell)/k$
    $\mathrm{BitwiseLT}_{k}(\Pi_{Bitwise-LT})$ $\mathrm{log}(\ell)$ $log(\ell)+2$ $k\ell\mathrm{log}(\ell)$ $\ell\mathrm{log}(\ell)+2\ell+2$
    $\mathrm{DReLU}_{k}(\Pi_{DReLU})$ $\mathrm{log}(\ell)+1$ $\mathrm{log}(\ell)+5$ $k(\ell\mathrm{log}(\ell)+4)$ $\ell\mathrm{log}(\ell)+2\ell+7+k$
    $\mathrm{ReLU}_{k}(\Pi_{ReLU})$ $\mathrm{log}(\ell)+2$ $\mathrm{log}(\ell)+6$ $k(\ell\mathrm{log}(\ell)+6)$ $\ell\mathrm{log}(\ell)+2\ell+9+k$
    $\mathrm{Maxpool}_{k}(\Pi_{Maxpool})$ $\mathrm{log}(k)(\mathrm{log}(\ell)+2)$ $\mathrm{log}(k)(\mathrm{log}(\ell)+6)$ $(k-1)(\ell\mathrm{log}(\ell)+6)$ $(1-1/k)(\ell\mathrm{log}(\ell)+2\ell+9+k)$
    TABLE II: The online running time and amortized communication costs of per party of the sub-protocol. Communication is in MB. Running time is in seconds. $n=21$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS. $\mathrm{Conv}_{32,32,64,5,1}$$\mathrm{MatMult}_{16,128,8192}$$\mathrm{ReLU}_{64}$$\mathrm{Maxpool}_{16,8,2}$.
    Operation t Protocol k Comm.(MB) WAN(s) LAN(s)
    $\mathrm{Conv}$ 3 LXY24 - 0.61 0.84 0.08
    Ours 8 0.09 0.15 0.05
    7 LXY24 - 0.80 0.87 0.30
    Ours 4 0.16 0.27 0.26
    $\mathrm{MatMult}$ 3 LXY24 - 1.30 1.84 0.19
    Ours 8 1.10 0.88 0.39
    7 LXY24 - 1.70 1.99 0.41
    Ours 4 1.16 0.92 0.41
    $\mathrm{ReLU}$ 3 LXY24 - 14.61 11.27 1.73
    Ours 8 5.31 7.42 1.40
    7 LXY24 - 16.96 12.62 1.55
    Ours 4 7.75 10.33 2.13
    $\mathrm{Maxpool}$ 3 LXY24 - 19.25 14.57 2.51
    Ours 8 7.00 10.09 2.06
    7 LXY24 - 22.35 16.86 3.85
    Ours 4 10.22 13.66 2.79
    VI-B Microbenchmarks
    We present the theoretical online round and communication complexity of the main protocols in Table I, and benchmark the online costs of the main sub-protocols for secure inference, including:
    • Matrix multiplication $\mathrm{MatMult}_{u,v,o}$ with truncation between a $u\times v$ matrix and a $v\times o$ matrix.
    • 2D convolution $(\mathrm{Conv}_{m,c,o,f,s})$ with the inputs of size $m\times m$, $c$ input channels, $o$ kernels of size $f\times f$ , and the stride $s$.
    • Secure ReLU $(\mathrm{ReLU}_{n})$ for a matrix of size $n\times n$.
    • Maxpool $(\mathrm{Maxpool}_{m,c,f})$ with $m\times m$ inputs, $c$ input channels, and $f\times f$ window.
    Concretely, we measure the running time and communication costs in both WAN and LAN settings with $n=21$ parties, where the size of $k$ is chosen randomly. The results are presented in Table II. The offline costs of the main sub-protocols are given in Appendix -A.
    For the convolution operations, we reduce the communication costs by $5.0-6.8\times$ compared to LXY24. The improvements arise from the use of our filter packing technique. Furthermore, we are $3.2-5.6\times$ faster in WAN and $1.15-1.6\times$ faster in LAN. We observe that the advantage increases as $k$ increases. For the matrix multiplication operations, we reduce the communication costs by $1.2-1.5\times$ compared to LXY24. This stems from our efficient vector-matrix multiplication. Moreover, we are $2.1\times$ faster in WAN. It should be noted that due to computational cost, our method offers little advantage in LAN. For the ReLU and Maxpool operations, we reduce the communication costs by $2.2-2.8\times$ and achieve $1.2-1.5\times$ speedups in the WAN setting, compared to LXY24. This benefits from our parallelizable non-linear protocol. However, in the LAN setting, the advantage becomes less significant.
@@ VI-B Microbenchmarks
    TABLE III: The amortized communication costs of per party of Liu et al.LXY24 and our protocol over different datasets and networks for secure inference. Communication is in MB. $n$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS, and OOM indicates memory overflow.
@@
    11 3 LXY24 - 71.77 40.45 112.22 105.63 59.53 165.16 165.04 99.40 264.44 1975.53 1113.65 3089.18
    4 LXY24 - 67.81 43.17 110.98 99.81 63.54 163.35 154.42 106.07 260.49 1866.63 1188.51 3055.14
    21 3 LXY24 - 78.73 36.97 115.70 115.89 54.42 170.31 183.03 90.86 273.89 2167.38 1017.94 3185.32
    7 LXY24 - 72.29 42.93 115.22 106.40 63.20 169.60 169.30 105.50 270.80 1989.81 1182.10 3171.91
    31 3 LXY24 - 80.89 35.62 116.51 119.07 52.42 171.49 188.64 87.54 276.18 2226.84 980.78 3207.62
    12 LXY24 - 70.30 44.85 115.15 103.47 66.02 169.49 159.74 110.21 269.95 OOM OOM OOM
    63 3 LXY24 - 83.06 34.11 117.17 122.26 50.21 172.47 194.29 83.85 278.14 OOM OOM OOM
    20 LXY24 - 75.26 42.84 118.10 110.78 63.05 173.83 172.48 105.27 277.75 OOM OOM OOM
    TABLE IV: Running time of Liu et al.LXY24 and our protocol over different datasets and networks for secure inference in LAN. Running time is in seconds. $n$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS, and OOM indicates memory overflow.
@@
    TABLE V: Running time of LXY24 and our protocol over different datasets and networks for secure inference in WAN. Running time is in seconds. $n$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS, and OOM indicates memory overflow.
    11 3 LXY24 - 17.30 9.64 26.94 23.80 13.92 37.72 35.35 22.15 57.50 426.93 182.71 609.64
    4 LXY24 - 16.82 9.93 26.75 23.51 13.21 36.72 34.34 22.43 56.77 420.59 183.26 603.85
    21 3 LXY24 - 32.82 15.68 48.50 48.14 22.02 70.16 64.87 35.80 100.67 914.20 307.68 1221.88
    7 LXY24 - 30.16 15.72 45.88 44.11 22.05 66.16 62.81 36.66 99.47 910.92 357.59 1268.51
    31 3 LXY24 - 48.14 20.09 68.23 67.59 28.47 96.06 95.06 46.27 141.33 1203.35 461.91 1665.26
    12 LXY24 - 44.12 22.01 66.13 63.73 30.32 94.05 90.15 49.95 140.10 OOM OOM OOM
    63 3 LXY24 - 89.99 36.69 126.68 131.76 55.08 186.84 188.09 86.27 274.36 OOM OOM OOM
    20 LXY24 - 87.73 45.19 132.92 124.20 64.03 188.23 180.91 106.59 287.50 OOM OOM OOM
    VI-C Benchmarks on Secure Inference